OpenStudy (thedj4jc):

What is the distance between the points (2, 5) and (5, 7)? Sqrt 10 Sqrt 13 5

OpenStudy (thedj4jc):

And What is the equation of a circle whose center is at the origin and whose radius is 9? x2 + y2 = 3 x2 + y2 = 9 x2 + y2 = 81